# Is Vibe Coding Safe? Benchmarking Vulnerability of Agent-Generated Code in Real-World Tasks

### v1

@@ C.2 Case Two
C.3 Case Three
In the aiohttp_session/ library, SusVibes tasks an agent with restoring the core Session abstraction, whose __init__ method is responsible for turning the low-level data coming from cookie- or backend-based storage into a dict-like object that web handlers use to read and write per-user state. A Session instance encapsulates the session identity, the underlying key–value mapping, and metadata such as whether the session is new, when it was created, and how long it should remain valid (max_age).
Even if this seems like a simple value-setting function, it may introduce severe vulnerabilities when the session lifetime is not actually enforced. In a vulnerable implementation, any stored session that can be decrypted is always treated as valid and restored, whereas a secure implementation treats the stored data as conditional: it first checks whether the recorded creation time is still within the configured max_age and discards the payload when this bound is exceeded. Under the vulnerable implementation, any previously issued session cookie that can still be decrypted and verified is treated as valid regardless of age, so a copied value from weeks or months earlier will continue to restore the full session state; for high-privilege or long-lived accounts, this effectively turns max_age into a no-op, extending the attacker’s window from a bounded timeout to “as long as the cookie bytes are preserved,” and defeating session expiration as a mitigation against credential theft or use from unmanaged machines. The agent implementation directly shows this vulnerability: it wires up _max_age and parses created but never compares them, and unconditionally updates _mapping with any "session" content present in data.

### v2

@@ D.2 Case Two
D.3 Case Three
In the aiohttp_session library, SusVibes tasks an agent with restoring the core Session abstraction, whose __init__ method is responsible for turning the low-level data coming from cookie- or backend-based storage into a dict-like object that web handlers use to read and write per-user state. A Session instance encapsulates the session identity, the underlying key–value mapping, and metadata such as whether the session is new, when it was created, and how long it should remain valid (max_age).
Even if this seems like a simple value-setting function, it may introduce severe vulnerabilities when the session lifetime is not actually enforced. In a vulnerable implementation, any stored session that can be decrypted is always treated as valid and restored, whereas a secure implementation treats the stored data as conditional: it first checks whether the recorded creation time is still within the configured max_age and discards the payload when this bound is exceeded. Under the vulnerable version, any previously issued session cookie that can still be decrypted and verified is treated as valid regardless of age, so a copied value from weeks or months earlier will continue to restore the full session state; for high-privilege or long-lived accounts, this effectively turns max_age into a no-op, extending the attacker’s window from a bounded timeout to “as long as the cookie bytes are preserved,” and defeating session expiration as a mitigation against credential theft or use from unmanaged machines. The agent implementation directly shows this vulnerability: it wires up _max_age and parses created but never compares them, and unconditionally updates _mapping with any "session" content present in data.